OpenStudy (anonymous):

How do you calculate the pH of acids & bases? Calculate pH 2.45 x 10-3 mol/L HCl? How do u calculate the 10-3 for example on a calculator? Or without one, i just dont get it

OpenStudy (anonymous):

Concentration of H+ = 2.45*10^-3. pH=-log(H+) = -log(2.45*10^-3)=3-log2.45 now calculate it(I assume you know how to work with logarithms if not That Logarithm is to base 10 so log(10^x)=x here x=-3 and log(ab)=loga+ log b here a=10^-3 and b=2.45)

OpenStudy (anonymous):

If you have a calculator, it needs to be a scientific calculator, with a LOG button on it somewhere for taking logarithms, and an EE or EXP button for entering numbers in scientific notation. If you have a smartphone, you can download a scientific calculator app from the Apple app store or the Android market for free. First, you need to enter the scientific notation. For the simplest scientific calculators, you enter the mantissa (the 2.45 part), then press an EE or EXP key and then enter the exponent (the -3 part). The result will show on the screen, in the simplest cases, as "2.45E-03" or sometimes it will look more like the regular scientific notation. Next, press the LOG key to take the common log (log base 10) of the number. Finally, press the sign change (+/-) key if the calculator has one. Or just remember to change the sign (+ to -, or - to +) of the answer, because pH(x) = - log(x), wihere x is your H+ concentration.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that is exactly what I needed, thank you so much. The scientific calculator on my iPhone is better than my actual scientific calculator. You're a lifesaver :)